OpenStudy (anonymous):

6 1/3 - 2 4/5 (show steps please)

OpenStudy (kropot72):

Turn the mixed numbers into fractions. 6 1/3=19/3 Can you do this to the second term?

OpenStudy (anonymous):

Yes.

OpenStudy (kropot72):

Gr8. So we have: 19/3 - 14/5. Do you follow?

OpenStudy (anonymous):

Yes.

OpenStudy (kropot72):

Next step is to combine the terms using a common denominator: ((19*5)-(14*3))/15 Make sense?

OpenStudy (anonymous):

Yes.

OpenStudy (kropot72):

So we get: (95 - 42)/15 = 53/15 = 3 8/15

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (kropot72):

You're welcome:)